An aqueous solution of NaOH having density 1.1kg/dm^3 contains 0.02 mole fraction of NaOH .What is the molality and molarity of NaOH solution?

Grade:11

1 Answers

Romani Ghosh
5 years ago
Molality of NaOH solution is no. of moles of NaOH dissolved in 1000g of water. 
No of moles of H2O = 1000/18 =55.55 mol 
Mole fraction of NaOH =0.02.
No of moles of NaOH = n2 /n2+n1 = n2/55.55 + n2 =0.02       so n2 = 1.134 mol
Molality of NaOH solution = 1.134 molal